# Ion counting and temperature determination of Coulomb-crystallized laser-cooled ions in traps using convolutional neural networks

### v1

@@ Abstract
Coulomb crystals, ordered structures of cold ions confined in ion traps, find applications in a variety of research fields. The number and temperature of the ions forming the Coulomb crystals are two key attributes of interest in many trapped-ion experiments. Here, we present a fast and accurate approach of determining these attributes from fluorescence images of the ions based on convolutional neural networks (CNNs). In this approach, we first generate a large number of images of Coulomb crystals with different ion numbers and temperatures using molecular-dynamics simulations and then train CNN models on these images to classify the desired attributes. The classification performance of several common pre-trained CNN models was compared in example tasks. We find that for crystals with ion numbers in the range 100–299 and secular temperatures of 5–15 mK, the best-performing model can discern number variations on the level of one ion with an accuracy of 93% and temperature variations by 1 mK with an accuracy of 92%. Since the trained model can be directly integrated into experiments, in-situ determination of these attributes can be realized in a non-invasive fashion, which has the potential to greatly facilitate the analysis and control of trapped ions in real time.
@@ I Introduction
Coulomb crystals, referring to the ordered structures of cold ions confined in ion traps, are an unusual form of matter that has been employed in a variety of research fields including quantum information and computation [1, 2, 3, 4], atomic clocks [5], fundamental physics studies [6] and cold and controlled chemistry[7]. Coulomb crystals can be generated by loading ions into radiofrequency (RF) or Penning traps [8] and cooling their motion to millikelvin or even lower temperatures using laser cooling (for suitable atomic ions) or sympathetic cooling (for atomic species that cannot be laser cooled and molecular ions). Since the ions are confined in deep traps in ultrahigh-vacuum chambers where collisions with gas and other perturbations from the surroundings are suppressed, Coulomb crystals can exhibit remarkable stability with lifetimes ranging from minutes to days, allowing the observation, addressing, and manipulation of the ions localized in the trap down to single particle level.
@@
For a Coulomb crystal, the ion number ($N$) and temperature ($T$) are two important attributes that are of interest in many experiments. For example, in studies of cold collisions between ions and other species, these two parameters and their variations can provide important information about the collision or reaction process in terms of collision energy and reaction rate [9]. During the past years, different methods have been developed to determine $N$ and $T$ of ions in Coulomb crystals. While $N$ is often obtained by counting ions with fluorescence imaging [10] and time-of-flight mass spectrometer (TOF-MS) [11, 12, 13, 14], the determination of $T$, often referred to as ion thermometry, is less straightforward and various methods have been developed applicable to different cases. For ions in Coulomb crystals with temperatures $T>1$ mK, thermometry techniques include the analysis of Doppler-broadened line shapes[15] and dark resonances [16], spatial imaging of single ions and small ion chains [17, 18, 19] as well as of large crystals[20, 21, 22, 23, 24]. For lower temperatures with ions cooled close to their motional ground state in the trap, resolved-sideband techniques [25, 26, 27, 28], applicable mainly to ion chains and small crystals, as well as extensions of this technique [29] applicable to large crystals, have been demonstrated in the past. For higher temperatures, Doppler-recooling [30, 31, 32, 33] and time-of-flight experiments [34] have been used.
@@
For Doppler-cooled Coulomb crystals, many of the ion counting and thermometry methods mentioned above rely on interference with the trapped ions, either by destroying the crystals, such as in the time-of-flight method where the ions are ejected from the trap, or potentially changing the ions’ attributes, as in Doppler-broadened line-shape analysis where the cooling-laser frequency is scanned and thus the ion temperatures changes. In the image-analysis methods, images of large crystals taken during experiments are compared visually to images generated by molecular dynamics (MD) simulations to infer $T$ and $N$ of the ions[20, 21, 22, 23, 24]. In the past, this visual comparison had to be performed a posteriori in a time-consuming, iterative simulation procedure and was not implemented in real-time which prevented a rapid and in-situ characterization of the Coulomb crystals during the experiments.
@@
Here, we present a fast, non-perturbative and accurate method to determine these two attributes based on convolutional neural networks (CNNs). In this approach, we first generated a large number of images of crystals of different $N$s and $T$s of ions in MD simulations and then trained CNN models on these images to classify these attributes. By comparing the classification results of several popular pre-trained CNN models, we identified the best models for determining $T$ and $N$, respectively. The trained model can be applied to the classification of experimental images and can also be integrated directly into experiments, making real-time and in-situ determination of crystals’ characteristics feasible, which offers perspectives to greatly facilitate the analysis and control of trapped ions in experiments.
@@
The paper is structured as follows. Sec. II explains the methods for generating a sufficient number of simulated images for training CNN models, the details of the selection of CNN models as well as the training and validation process. Sec. III provides some examples of applying CNN models to image classification tasks, compares the performances of several CNN models, as well as discusses the implementation of the trained models to classify experimental results. Sec. IV summarizes the main results and suggests potential avenues for further improvements.
@@ II.1 Image generation
Figure 1: (a) An example false-color image of a Coulomb crystal with 200 ions and a secular temperature 10 mK. (b) Fluctuation of the temperature after stabilization extracted from the simulation generating the crystal in (a).
Figure 2: (a) Example false-color images of Coulomb crystals with different numbers of ions for a fixed secular temperature of $T=10$ mK. (b) Example images of Coulomb crystals with different secular temperatures for a fixed number of ions $N=200$.
For training a CNN model, the quality and quantity of images input into the model are crucial for its ultimate performance in terms of accuracy and reliability. Although experimental images of Coulomb crystals taken by cameras would be ideal sources for this task, the collection of such a vast amount of data is painstaking. To our knowledge, to date there is no systematic collection with a sufficient number of images covering a wide range of well-determined values for the salient attributes available. Since image comparison between simulations and experiments is employed routinely as a method for determining $N$ and $T$, generating a large number of images with varying attributes by simulation was pursued as an efficient way to obtain suitable images for training the CNN model, which was then transferred to the classification of experimental images.
To generate images of Coulomb crystals with a well-defined number and temperature of ions, we adapted the molecular-dynamics (MD) simulation method described in [35]. Here, the trajectories of $N$ ions inside a RF ion trap were simulated by taking into account the different forces acting on the trapped ions. The force $\bm{F}_{\rm{i}}$ exerted on an ion $i$ can be written as
@@
where $U(\bm{r},t)$ is the time-dependent trapping potential, $\bm{F}_{\rm{Coulomb}}$ the Coulomb interaction between ions, $\bm{F}_{\rm{cooling}}$ the cooling force from photon scattering, and $\bm{F}_{\rm{bg}}$ an effective heating force assumed to be dominated by collisions with background gas.
Generally, the trap potential $U(\bm{r},t)$ is determined by the trap geometry and experimental parameters (DC voltages and RF amplitude/frequency applied to different electrodes). The specific implementation of the trap varies from one experiment to another. In the present work, we used the harmonically approximated potential of a segmented RF trap described in [36],
@@
where $m$ is the mass of the ion, $\omega$ the applied RF frequency, $r_{i}$ the ion position coordinates. Here, $\Omega$ was set to $2\pi\times 10$ MHz, and the Mathieu parameters $a_{i}$ and $q_{i}$ [8] were derived from numerical modeling: $a_{x}=-4.62\times 10^{-4}$, $a_{y}=-4.10\times 10^{-4}$, $a_{z}=8.83\times 10^{-4}$, $q_{x}=8.73\times 10^{-2}$, $q_{y}=-9.15\times 10^{-2}$, and $q_{z}=0$ thus mimicking realistic experimental values.
MD simulations were carried out using the OpenMM framework [37], a high-performance, open-source toolkit supporting both built-in force fields such as Coulomb interactions as well as custom forces and allowing for a speed-up of the computations using graphics processing unit (GPU) acceleration. The cooling force $\bm{F}_{\rm{cooling}}$ was modeled by momentum kicks of magnitude $\hbar{k}$ due to photon absorption or emission applied to the ions as they undergo transitions between ground and excited states during laser cooling [35] .
While the number of ions $N$ can simply be specified as an input parameter for the MD simulation, the temperature of ions $T$ is less straightforward to control as it depends on the dynamics of the ion ensemble in the presence of several competing interaction terms. One common way to regulate $T$ is to repeatedly apply a velocity kick with a random direction to the ions which represents an overall effective heating effect for various heating processes [22]. However, as examined in [35], adding a velocity kick which essentially resembles an elastic collision of ions with other species (or virtual particles [23]) can cause the ions to sample a broad range of temperatures, especially for infrequent kicks with a large momentum transfer. According to our tests, even adding continuous, weak velocity kicks can cause two severe problems for generating images with a well-defined $T$. First, the stability of $T$ during the simulation can easily be worse than $\pm$1 mK, which would lead to images difficult for the CNN model to classify with an accuracy on that level. Second, there is not a universal velocity-kick frequency and magnitude to generate images with a certain $T$ if their $N$’s are different, which means a tedious trial-and-error procedure is needed to tune the kick parameters to obtain a desired $T$ once an image with a different $N$ is desired.
@@
where $\overline{v}_{ix}$, $\overline{v}_{iy}$, and $\overline{v}_{iz}$ are the velocity components of the $i$-th ion averaged over one RF period. In this way, the contribution of the micromotion [8] to the ion velocities was removed from the averaging resulting in a temperature that purely reflects the secular, i.e., thermal, motion of the ions in the trap. The difference between $T_{\rm{period}}$ and $T_{\rm{target}}$ was then employed as an error signal $e(t)$ to modify the velocities of the ions based on a PID controller. The modified velocity $u_{i}$ of the $i$-th ion was related to its previous velocity $v_{i}$ by
@@
where $K_{P}$, $K_{I}$, $K_{D}$ are the coefficients for the proportional (P), integral (I), and derivative (D) terms in the controller. In practice, we found that only using the P term was sufficient for the stabilization. We then defined an average of $T_{\rm{period}}$ over 100 RF periods as the secular temperature of the ion ensemble $T$.
After $T$ was stabilized, the position of each ion after each time step (1 ns) was recorded and binned into a 3D histogram of positions collected over a duration of 1 ms. From the 3D histogram, 2D slices were extracted along the viewing direction of a camera. For each slice, a Gaussian blur was applied, the intensity of which was varied based on the slice’s distance from the central layer to simulate the effect of the finite focal depth of the microscope attached to the camera. All slices were then combined, normalized, and saved as a final simulated image. As an example, Fig. 1 (a) shows an image of a Coulomb crystal with $T$=10 mK and $N$=200 thus generated and Fig.1(b) the corresponding time-dependent temperature after stabilization. It can be seen that the fluctuation of $T$ is within $\pm$0.04 mK over 1 ms owing to the PID stabilization. This amounts to $\pm$4relative uncertainty with respect to 1 mK, which is sufficiently precise to serve as training data for a CNN model.
In this way, a large number of simulated images with varying $N$ and $T$ was generated to train the models. As an example, Fig. 2(a) shows a small collection of images of Coulomb crystals generated with different numbers of ions at a secular temperature of 10 mK, and Fig. 2(b) displays crystals with 200 ions at different secular temperatures.
@@ II.2 CNN models
To classify the images of Coulomb crystals in terms of $N$ and $T$, we employed one of the most widespread deep-learning methods for image classification, i.e.,, CNNs[38, 39]. CNNs can automatically and adaptively learn spatial hierarchies of features from input images by applying convolutional and pooling layers, and have proven to be especially effective in image recognition [39, 40, 41, 42].
As a common practice when using CNNs, we adopted transfer learning [43, 44, 45] for the present image-classification task. This technique reuses a pre-trained model from other tasks as the starting point for a new task, rather than starting from scratch. During transfer learning, the initial layers of the CNN models can capture generic features while the later layers are fine-tuned to the new task. This approach leverages the knowledge gained from the initial training to accelerate and enhance the new learning process, thus leading to faster convergence and better performance.
Among a large number of pre-trained CNN models that are available for transfer learning, we selected a few widely-used variants, i.e., AlexNet[39], ResNet18[41], Vgg16[40], and MnasNet[42], for our task in consideration of a balance between their performance and computational efficiency. The definitions of these models, including their architectures and the pre-trained weights, are accessible in the $torchvision.models$ module of PyTorch [46], an open-source machine learning framework that was utilized in the present work.
@@ II.3 Training and validation
To use the generated images for training a CNN model, the range of $N$ (or $T$) to be classified by the model needs to be specified. For example, a given range of $N=100--299$ with a step of 1 will train the model to classify the number of ions in the crystals to differences within one ion count. The specified images are resized to 224$\times$224 pixels as required by the pre-trained CNN models mentioned above. In addition, the original grayscale images can be converted to RGB images. Interestingly, in the present classification it turned out that grayscale images outperformed RGB images for classifying $N$ whereas RGB images gave better accuracy for classifying $T$. This difference could be explained by noting that for classifying $N$, shapes and sizes of crystals are more important features in comparison with colors, whereas for classifying $T$, the additional two color channels of RGB images could offer additional dimensionality for feature extraction, enabling the model to capture more complex patterns across channels.
@@
As we will discuss below, $N$ could be reliably determined by the present models with high accuracy for a much larger number of classes than $T$. Although it is possible to train a model that can simultaneously classify both $N$ and $T$, known as multi-label classification, we found that the classification accuracy was significantly lower compared to training models for $N$ or $T$ separately. Therefore, in the remainder of the article, we will focus on the training of individual models.
@@
As an example of classifying $N$=100–299 (in steps of 1 ion), we prepared a total of 7200 images, i.e., 36 images for each class of $N$, and divided them randomly into training and validation datasets at a ratio of 4:1. A batch size of 32, chosen based on our GPU capacity, is used for both training and validation. All the CNN models mentioned above can be employed either with or without its pre-trained weights (PTW); the latter means that only the architecture of the model is used. A comparison of classification performances of different models with or without using PTW will be given in the next section, with the help of specific image-classification examples.
@@
After loading the model, its first convolutional layer was modified to accept single-channel input if the images were grayscale, as all the models by default accept three channels for RGB images. Morevoer, the last layer of the model was also modified according to the number of classes to be classified. In our example, the last layer was adapted to 200 classes for classifying $N$=100–299 with a step of 1. We used the cross-entropy loss function [39, 47] to measure the dissimilarity between the predicted probability distribution and the true distribution of the target classes, and the Stochastic Gradient Descent (SGD) optimizer with a learning rate of 0.0006 and momentum of 0.9 to update the CNN parameters during the training loop [47].

### v2

@@ Abstract
Coulomb crystals—ordered structures of cold ions confined in ion traps—find applications in a variety of research fields. The number and temperature of the ions forming the Coulomb crystals are two key attributes of interest in many trapped-ion experiments. Here, we present a fast and accurate approach to determining these attributes from fluorescence images of the ions based on convolutional neural networks (CNNs). In this approach, we first generate a large number of images of Coulomb crystals with different ion numbers and temperatures using molecular-dynamics simulations and then train CNN models on these images to classify the desired attributes. The classification performance of several common pretrained CNN models was compared in example tasks. We find that for crystals with ion numbers in the range 100–299 and secular temperatures of 5–15 mK, the best-performing model can discern number variations on the level of one ion with an accuracy of 93% and temperature variations by 1 mK with an accuracy of 92%. Since the trained model can be directly integrated into experiments, in-situ determination of these attributes can be realized in a non-invasive fashion, which has the potential to greatly facilitate the analysis and control of trapped ions in real time.
@@ I Introduction
Coulomb crystals—ordered structures of cold ions confined in ion traps—are an unusual form of matter that has been employed in a variety of research fields including quantum information and computation [1, 2, 3, 4], atomic clocks [5], fundamental physics studies [6] and cold and controlled chemistry[7]. Coulomb crystals can be generated by loading ions into radio-frequency (RF) or Penning traps [8] and cooling their motion to millikelvin or even lower temperatures using laser cooling (for suitable atomic ions) or sympathetic cooling (for atomic species that cannot be laser cooled and molecular ions). Since the ions are confined in deep traps in ultrahigh-vacuum chambers where collisions with gas and other perturbations from the surroundings are suppressed, Coulomb crystals can exhibit remarkable stability with lifetimes ranging from minutes to days, allowing the observation, addressing, and manipulation of the ions localized in the trap down to the single- particle level.
@@
For a Coulomb crystal, the number of ions ($N$) and temperature ($T$) are two important attributes that are of interest in many experiments. For example, in studies of cold collisions between ions and other species, these two parameters and their variations can provide important information about the collision or reaction process in terms of collision energy and reaction rate [9]. During the past years, different methods have been developed to determine $N$ and $T$ for ions in Coulomb crystals. While $N$ is often obtained by counting ions using fluorescence imaging [10] or a time-of-flight mass spectrometer (TOF-MS) [11, 12, 13, 14], the determination of $T$, often referred to as ion thermometry, is less straightforward, and various methods have been developed that are applicable to different cases. For ions in Coulomb crystals with temperatures $T>1$ mK, thermometry techniques include the analysis of Doppler-broadened line shapes[15] and dark resonances [16], spatial imaging of single ions and small ion chains [17, 18, 19] as well as of large crystals[20, 21, 22, 23, 24]. For lower temperatures with ions cooled close to their motional ground state in the trap, resolved-sideband techniques [25, 26, 27, 28], applicable mainly to ion chains and small crystals, as well as extensions of this technique [29] applicable to large crystals, have been demonstrated in the past. For higher temperatures, Doppler-recooling [30, 31, 32, 33] and time-of-flight experiments [34] have been used.
@@
For Doppler-cooled Coulomb crystals, many of the ion-counting and thermometry methods mentioned above rely on interference with the trapped ions, either by destroying the crystals, such as in the time-of-flight method, where the ions are ejected from the trap, or potentially changing the ions’ attributes, as in Doppler-broadened line-shape analysis, where the cooling-laser frequency is scanned and thus the ion temperature changes. In the image-analysis methods, images of large crystals taken during experiments are compared visually to images generated by molecular-dynamics (MD) simulations to infer $T$ and $N$ for the ions[20, 21, 22, 23, 24]. In the past, this visual comparison had to be performed a posteriori in a time-consuming, iterative simulation procedure and was not implemented in real-time, which prevented a rapid and in-situ characterization of the Coulomb crystals during experiments.
@@
Here, we present a fast, nonperturbative, and accurate method to determine these two attributes based on convolutional neural networks (CNNs). In this approach, we first generated a large number of images of crystals with different $N$ and $T$ values using MD simulations and then trained CNN models on these images to classify these attributes. By comparing the classification results of several popular pretrained CNN models, we identified the best models for determining $T$ and $N$, respectively. The trained model can be applied to the classification of experimental images and can also be integrated directly into experiments, making real-time and in-situ determination of crystals’ characteristics feasible, which offers the potential to greatly enhance the analysis and control of trapped ions in experiments.
@@
The remainder of this paper is structured as follows. Sec. II explains the methods for generating a sufficient number of simulated images for training CNN models, and it presents the details of the selection of CNN models as well as the training and validation process. Sec. III provides some examples of the application of CNN models to image-classification tasks, compares the performances of several CNN models, and discusses the application of the trained models to classifying experimental results. Sec. IV summarizes the main results and suggests potential avenues for further improvements.
@@ II.1 Image generation
Figure 1: (a) An example false-color image of a Coulomb crystal with 200 ions and a secular temperature of 10 mK. (b) Fluctuation of the temperature after stabilization extracted from the simulation generating the crystal in (a).
Figure 2: (a) Example false-color images of Coulomb crystals with different numbers of ions for a fixed secular temperature of $T=10$ mK. (b) Example images of Coulomb crystals with different secular temperatures for a fixed number of ions $N=200$. The color bar applies across all figures showing crystal images in the article.
For training a CNN model, both the quality and quantity of images input into the model are crucial for its ultimate performance in terms of accuracy and reliability. Although experimental images of Coulomb crystals taken by cameras would be ideal sources for this task, the collection of such a vast amount of data would be painstaking. To our knowledge, there is currently no systematic collection with a sufficient number of images covering a wide range of well-determined values for the salient attributes available. Since image comparison between simulations and experiments is routinely employed as a method for determining $N$ and $T$, generating a large number of images with varying attributes by simulation was pursued as an efficient way to obtain suitable images for training the CNN model, which was then transferred to the classification of experimental images.
To generate images of Coulomb crystals with well-defined numbers and temperatures of ions, we adapted the MD simulation method described in Ref. [35]. Here, the trajectories of $N$ ions inside an RF ion trap were simulated by taking into account the different forces acting on the trapped ions. The force $\bm{F}_{\rm{i}}$ exerted on an ion $i$ can be written as
@@
where $U(\bm{r},t)$ is the time-dependent trapping potential, $\bm{F}_{\rm{Coulomb}}$ is the Coulomb interaction between ions, $\bm{F}_{\rm{cooling}}$ is the cooling force from photon scattering, and $\bm{F}_{\rm{bg}}$ is an effective heating force assumed to be dominated by collisions with background gas, which also effectively includes other heating mechanisms such as RF heating [35].
Generally, the trap potential $U(\bm{r},t)$ is determined by the trap geometry and experimental parameters (DC voltages and the RF amplitude and frequency applied to different electrodes). The specific implementation of the trap varies from one experiment to another. In the present work, we used the harmonically approximated potential of a segmented RF trap, under the experimental conditions described in detail in Ref. [36]. The potential takes the form
@@
where $m$ is the mass of the ion, $\omega$ is the applied RF frequency, and $r_{i}$ is the ion position coordinates. Here, $\Omega$ was set to $2\pi\times 10$ MHz, and the Mathieu parameters $a_{i}$ and $q_{i}$ [8] were derived from numerical modeling: $a_{x}=-4.62\times 10^{-4}$, $a_{y}=-4.10\times 10^{-4}$, $a_{z}=8.83\times 10^{-4}$, $q_{x}=8.73\times 10^{-2}$, $q_{y}=-9.15\times 10^{-2}$, and $q_{z}=0$ thus mimicking realistic experimental values.
MD simulations were carried out using the OpenMM framework [37], a high-performance, open-source toolkit supporting both built-in force fields, such as Coulomb interactions, and custom forces. This framework also enables the computations to be sped up using graphics-processing-unit (GPU) acceleration. The cooling force $\bm{F}_{\rm{cooling}}$ was modeled by momentum kicks of magnitude $\hbar{k}$ due to photon absorption or emission applied to the ions as they undergo transitions between ground and excited states during laser cooling [35] .
While the number of ions $N$ can simply be specified as an input parameter for the MD simulations, the temperature of the ions $T$ is less straightforward to control, as it depends on the dynamics of the ion ensemble in the presence of several competing interaction terms. One common way to regulate $T$ is to repeatedly apply a velocity kick with a random direction to the ions, which represents an overall effective heating effect for various heating processes [22]. However, as examined in Ref. [35], adding a velocity kick that essentially resembles an elastic collision of ions with other species (or virtual particles [23]) can cause the ions to sample a broad range of temperatures, especially for infrequent kicks with a large momentum transfer. According to our tests, even adding continuous, weak velocity kicks can cause two severe problems for generating images with a well-defined $T$. First, the stability of $T$ during the simulation can easily be worse than $\pm$1 mK, which would lead to images that are difficult for the CNN model to classify with an accuracy on that level. Second, there is not a universal velocity-kick frequency and magnitude to generate images with a certain $T$ if their $N$ values are different, which means that a tedious trial-and-error procedure is needed to tune the kick parameters to obtain a desired $T$ once an image with a different $N$ is desired.
@@
where $\overline{v}_{ix}$, $\overline{v}_{iy}$, and $\overline{v}_{iz}$ are the velocity components of the $i$th ion averaged over one RF period. In this way, the contribution of the micromotion [8] to the ion velocities was removed from the averaging, resulting in a temperature that purely reflects the secular, i.e., thermal, motion of the ions in the trap. The difference between $T_{\rm{period}}$ and $T_{\rm{target}}$ was then employed as an error signal $e(t)$ to modify the velocities of the ions based on a proportional–integral–derivative (PID) controller. The modified velocity $u_{i}$ of the $i$th ion was related to its previous velocity $v_{i}$ by
@@
where $K_{P}$, $K_{I}$, and $K_{D}$ are the coefficients for the proportional (P), integral (I), and derivative (D) terms in the controller. In practice, we found that using only the P term was sufficient for stabilization. The optimized parameters used in our simulations are $K_{P}=0.65$ and $K_{I}=K_{D}=0$. We then defined an average of $T_{\rm{period}}$ over 100 RF periods as the secular temperature of the ion ensemble $T$.
After $T$ was stabilized, the position of each ion after each time step (1 ns) was recorded and binned into a three-dimensional (3D) histogram of positions collected over a duration of 1 ms. From the 3D histogram, two-dimensional slices were extracted along the viewing direction of a camera. For each slice, a Gaussian blur was applied, the intensity of which was varied based on the slice’s distance from the central layer to simulate the effect of the finite focal depth of the microscope attached to the camera. All slices were then combined, and the resulting image was normalized to its maximum intensity before being saved as the final simulated image. As an example, Fig. 1 (a) shows an image of a Coulomb crystal with TT=10 mK and NN=200 thus generated, and Fig.1(b) shows the corresponding time-dependent temperature after stabilization. It can be seen that the fluctuation of $T$ is within $\pm$0.04 mK over 1 ms due to the PID stabilization. This amounts to $\pm$4% relative uncertainty with respect to 1 mK, which is sufficiently precise to serve as training data for a CNN model. In comparison to other approaches to simulating crystals relying on manual parameter tuning, our PID-based method offers automated and more precise temperature control, enabling efficient generation of high-quality training data, which is essential for training robust models.
In this way, a large number of simulated images with varying $N$ and $T$ were generated to train the models. As an example, Fig. 2(a) shows a small collection of images of Coulomb crystals generated with different numbers of ions at a secular temperature of 10 mK, and Fig. 2(b) displays crystals with 200 ions at different secular temperatures.
@@ II.2 CNN models
To classify the images of Coulomb crystals in terms of $N$ and $T$, we employed one of the most widespread deep-learning methods for image classification: CNNs [38, 39]. CNNs can automatically and adaptively learn spatial hierarchies of features from input images by applying convolutional and pooling layers, and they have proven to be especially effective in image recognition [39, 40, 41, 42].
As a common practice when using CNNs, we adopted transfer learning [43, 44, 45] for the present image-classification task. This technique reuses a pretrained model from other tasks as the starting point for a new task, rather than starting from scratch. During transfer learning, the initial layers of the CNN models can capture generic features while the later layers are fine-tuned to the new task. This approach leverages the knowledge gained from the initial training to accelerate and enhance the new learning process, thus leading to faster convergence and better performance.
Among a large number of pretrained CNN models that are available for transfer learning, we selected a few widely used variants—AlexNet [39], ResNet18 [41], VGG16 [40], and MnasNet [42]—for our task in consideration of a balance between their performance and computational efficiency. The definitions of these models, including their architectures and the pretrained weights, are accessible in the torchvision.models module of PyTorch [46], an open-source machine-learning framework that was used in the present work. For example, we can specify the model architecture and load the pretrained weights by calling
    from torchvision import models
    net = models.resnet18(pretrained=True)
This loads the ResNet18 model with weights pretrained on the ImageNet dataset [47], a large-scale benchmark dataset containing over one million labeled images across 1000 object categories.
@@ II.3 Training and validation
To use the generated images for training a CNN model, the range of $N$ (or $T$) to be classified by the model needs to be specified. For example, a given range of NN=100–299 with a step of 1 will train the model to classify the number of ions in the crystals to differences within one ion count. The specified images are resized to $\times224$ pixels as required by the pretrained CNN models mentioned above. In addition, the original grayscale images can be converted to Red–Green–Blue (RGB) images using the Image.convert(‘RGB’) method provided by the Python Imaging Library [48], which replicates the grayscale channel across the three RGB channels. This transformation does not introduce new information but ensures compatibility with CNNs pretrained on RGB images. Interestingly, in the present classification, it turned out that grayscale images outperformed RGB images for classifying $N$, whereas RGB images gave better accuracy for classifying $T$. This difference could be explained by noting that for classifying $N$, the shapes and sizes of crystals are more important features than colors, whereas for classifying $T$, the additional two color channels of RGB images could offer additional dimensionality for feature extraction, enabling the model to capture more complex patterns across channels.
@@
As we will discuss below, $N$ could be reliably determined by the present models with high accuracy for a much larger number of classes than $T$. Although it is possible to train a model that can simultaneously classify both $N$ and $T$, known as multilabel classification, we found that the classification accuracy was significantly lower compared to training models for $N$ or $T$ separately. Therefore, in the remainder of the article, we will focus on the training of individual models.
@@
As an example of classifying NN=100–299 (in steps of 1 ion), we prepared a total of 7200 images, i.e., 36 images for each class of $N$, and we divided them randomly into training and validation datasets in the ratio 4:1. A batch size of 32, chosen based on our GPU capacity, was used for both training and validation. All the CNN models mentioned above can be employed either with or without their pretrained weights (PTWs); the latter means that only the architecture of the model is used. A comparison of the classification performances of different models with or without using PTWs will be given in the next section, with the help of specific image-classification examples.
@@
After loading the model, its first convolutional layer was modified to accept single-channel input if the images were grayscale, as all the models by default accept three channels for RGB images. Moreover, the last layer of the model was also modified according to the number of classes. In our example, the last layer was adapted to 200 classes for classifying NN=100–299 with a step of 1. We used the cross-entropy loss function [39, 49] to measure the dissimilarity between the predicted probability distribution and the true distribution of the target classes, and the stochastic-gradient-descent (SGD) optimizer with a learning rate of 0.0006 and momentum of 0.9 to update the CNN parameters during the training loop [49].